# Mind the GAP: Glimpse-based Active Perception improves generalization and sample efficiency of visual reasoning

### v1

@@ 3 Related work
Inspired by human active vision, several models for learning visual relations have been introduced. By decomposing input images into a sequence of image regions obtained from synthetic eye movements, Woźniak et al. (2023) demonstrated that visual reasoning tasks can be solved with high accuracy using relatively compact networks. However, the synthetic eye movements were based on hard-coded locations, whereas in our case, the glimpsing process is controlled by a saliency map. Aiming primarily at computationally efficient image classification, Mnih et al. (2014) used reinforcement learning to train a glimpsing policy. The policy iteratively determines the next 2D glimpse location based on the history of previously visited locations and the corresponding glimpse contents. However, having been evaluated on several visual reasoning tasks by Vaishnav & Serre (2023), this model could not achieve reasonable performance. This is likely caused by learning inefficiency in exploring the space of all possible 2D glimpse locations. Our approach, in contrast, leverages the concept of salience maps to determine the next glimpse location which significantly reduces the space of glimpse locations to the most salient ones. Gregor et al. (2015) proposed a variational auto-encoder that uses an attention mechanism for an iterative (re-)construction of complex images by attending to their different parts. Its extension proposed by Adeli et al. (2023) was designed to tackle visual reasoning tasks where it outperformed baseline CNNs. However, being trained by image reconstruction, the model does not always generalize well to OOD visual inputs. Vaishnav & Serre (2023) introduced a recurrent soft attention mechanism guided by a memory module that masks specific parts of the image where the processing should be focused. At each iteration, the model can thereby attend to multiple regions of the image simultaneously which is different from our approach where only one region can be focused at a time.
@@
Another family of models combines object-centric representations (Greff et al., 2019; Locatello et al., 2020) with relational inductive biases (Webb et al., 2021; Kerg et al., 2022; Webb et al., 2024b). The former decomposes an image into objects present in it while the latter constrains the processing to relations between objects instead of their specific properties. Webb et al. (2024a) used pre-trained slot attention (Locatello et al., 2020) to segment objects into a sequence of slots that contain disentangled representations of objects’ visual features and their positional information. Each pair of slots is then processed by a relational operator and the resulting sequence of relational representations is processed by the Transformer (Vaswani et al., 2017). The successor model, Slot-Abstractor (Mondal et al., 2024), processes the sequence of slots with a recently introduced Abstractor module (Altabaa et al., 2024) that proved to be especially effective in reasoning about relations between objects using their disentangled visual and positional features. Similar to our approach, there are two streams of information: one based on visual features and one based on positional information. However, the latter stream consists of positional embeddings that are learned during the pre-training process which may make the model more fragile to OOD inputs. In contrast, our approach uses the raw 2D locations of salient regions.
@@ 4 Experiments
Our approach was evaluated on four visual reasoning datasets illustrated in Figure 3. We start with an evaluation on all 23 tasks of the SVRT dataset (Fleuret et al., 2011) to see how efficiently the model can learn spatial and similarity relations between simple shapes, see an example in Figure 3(a). Next, we test the model’s OOD capabilities to generalize learned relations to previously unseen shapes. Specifically, we use an extended version of the task #1 from the SVRT dataset (Puebla & Bowers, 2022), referred to as SVRT #1-OOD, see Figure 3(b). It contains test sets with OOD shapes coming from 13 different distributions. Additionally, we use the ART dataset (Webb et al., 2021), which contains more abstract relations, see Figure 3(c). Finally, we test the model’s potential to scale to more complex images by testing it on the CLEVR-ART dataset (Webb et al., 2024a), see Figure 3(d). It contains two tasks from ART but with more realistic images. Each dataset is described in more detail in Appendix B.
@@ 5 Results
In all experiments, we report the performance of GAP combined with two downstream architectures, Transformer and Abstractor. The corresponding models are referred to as GAP-Transformer and GAP-Abstractor respectively. Details about hyper-parameters, such as the type of glimpse sensor (multi-scale or log-polar), and training process are described in Appendix C. Moreover, extensive performance reports for different glimpse sensors are included in Appendix  D.
Table 1: Test accuracy for SVRT. The results obtained from the best trained models are averaged over SD, SR and over all (SD + SR) tasks. The standard deviation is calculated over accuracies for each individual task. Prior art results are listed in the top section whereas results obtained from our evaluations are listed in the two bottom sections. The highest and the second-highest mean accuracy in each column is marked in bold and underlined, respectively. The models appended by * contain a module pre-trained on the train sets from all SVRT tasks to segment the shapes.
@@
Figure 4: Sample efficiency evaluation for SVRT. Average test accuracy on all 23 SVRT tasks depending on the size of the training dataset. Panel (a) compares our best model with prior art. Panel (b) compares our glimpse-based models with their ablated counterparts.
@@ 5.2 Out-of-distribution generalization of same-different relation
While the SVRT dataset allows us to see whether a model can learn visual relations from a limited amount of training data, it does not allow us to test the model’s ability to generalize those relations to OOD inputs. We evaluate the OOD generalization of a single same-different relation using the SVRT #1-OOD dataset (Appendix  B.2) where a model has to determine whether two shapes are the same or different up to translation. Importantly, the models are evaluated on test sets with OOD shapes coming from 13 different distributions (see Figure B.2 for examples). Table 5.2 shows the performance averaged over all OOD test sets and Figure 5 shows the performance on each of the test sets. Outperforming the prior art models, our model, GAP-Abstractor, performs quite consistently on all test sets being slightly less accurate on only two sets (’straight_lines’ and ’rectangles’).
@@ 5.4 Scaling potential for more complex visual scenes
We test our approach on two tasks from ART (RMTS and ID) but with more realistic images using the CLEVR-ART dataset. While the images from other datasets considered in this work are binary and contain only 2D objects, the CLEVR-ART images are colored and contain 3D objects with such real-world features as shades and reflections. Hence, the glimpsing behavior can be distracted by spurious salient locations such as edges of shades or regions of high reflection. Table 4 highlights that the GAP-Abstractor performs on par with the state-of-the-art model, Slot-Abstractor, being marginally better in one task and marginally worse in the other. However, Slot-Abstractor was pre-trained on all shapes from CLEVR-ART being thereby potentially exposed to objects from both train and test sets. This is in contrast to our models, which were not pre-trained and were exposed only to the objects from the train set. To compare our models with Slot-Abstractor on a more equal footing, we generated two additional OOD test sets shown in Figure 6. The new OOD-1 test set contains one novel type of shape – pyramid – on which the Slot-Abstractor was not pre-trained. Similarly, the new OOD-2 test set contains 2 such novel shapes: pyramids and toruses. The corresponding results are shown in Figure 7 demonstrating superior OOD generalization of our GAP-based model compared with Slot-Abstractor. Notably, the latter performs significantly worse, if the images contain multiple objects on which it was not pre-trained.
@@ 6 Discussion
Our results suggest that factorizing an image into its complementary “what” and “where” contents plays an essential role. The low-dimensional “where” content (glimpse locations) is crucial since it does not contain any visual information and, in turn, allows to learn relations that generalize to OOD images. To capitalize on that, it is important to process the “where” content explicitly, disentangling it from its “what” complement, so that the latter carries only a modulatory effect in form of attention. This is achieved by employing the recently introduced Abstractor as our downstream architecture. In contrast, implicitly mixing the “what” and the “where” contents, as in the case of the Transformer downstream architecture, weakens the generalization capabilities. This can be seen by comparing the performance between GAP-Transformer and GAP-Abstractor. To further support this, we provide supplementary results in Appendix D.1 showing the importance of using both the “what” and the “where” contents instead of just one of them for task solving.
Another important aspect of our model is that it processes only the most salient image regions while ignoring the unimportant ones. The inferior performance of models where the downstream architectures receive full information (i.e. all patches that constitute the image) suggests that supplying a superfluous amount of information may distract the models hindering the learning process. Additional results provided in Appendix D.2 elucidate this further. Specifically, we show that it is insufficient to naively reduce the amount of information provided to the downstream architecture, by discarding uninformative image parts. Instead, it has to be ensured that the supplied information is well structured. In the context of GAP, it means that the glimpse contents have to be centered around salient image regions instead of, for example, being picked from a set of non-overlapping patches obtained by splitting the image using a regular grid (as is done in ViT-based models).
@@ Appendix D Ablations and additional results
GAP-Transformer mult-scale 76.4 ± 4.3 64.7 ± 5.6
@@
Dataset size SD tasks
@@
D.1 Evaluating importance of glimpse contents and glimpse locations
@@
D.2 Evaluating the role of processing only the relevant information

### v2

@@ 3 Related work
Inspired by human active vision, several models for learning visual relations have been introduced. By decomposing input images into a sequence of image regions obtained from synthetic eye movements, Woźniak et al. (2023) demonstrated that visual reasoning tasks can be solved with high accuracy using relatively compact networks. However, the synthetic eye movements were based on hard-coded locations, whereas in our case, the glimpsing process is controlled by a saliency map. The approach from (Larochelle & Hinton, 2010) and its ANN-based extension (Mnih et al., 2014) use reinforcement learning (RL) to train a glimpsing policy to determine the next glimpse location either in a continuous 2D space or in a discrete grid-based space of all possible glimpse locations. While this approach was evaluated mainly on simple image classification and object detection tasks, Ba et al. (2014); Xu et al. (2015) extended it to more complex images and image captioning tasks. However, having been evaluated on several visual reasoning tasks by Vaishnav & Serre (2023), the RL-based approaches could not achieve reasonable performance. This is likely caused by learning inefficiency in exploring the large space of all possible glimpse locations. Nevertheless, the RL-based approaches that use more efficient RL techniques such as (Pardyl et al., 2025) to learn complex glimpsing policies are relevant to our work as they can be integrated into our model to enhance its capabilities of dealing with real-world images. Our approach, in contrast, leverages the concept of saliency maps to determine the next glimpse location which significantly reduces the space of glimpse locations to the most salient ones. Gregor et al. (2015) proposed a variational auto-encoder that uses an attention mechanism for an iterative (re-)construction of complex images by attending to their different parts. Its extension proposed by Adeli et al. (2023) was designed to tackle visual reasoning tasks where it outperformed baseline CNNs. However, being trained by image reconstruction, the model does not always generalize well to OOD visual inputs. Vaishnav & Serre (2023) introduced a recurrent soft attention mechanism guided by a memory module that masks specific parts of the image where the processing should be focused. At each iteration, the model can thereby attend to multiple regions of the image simultaneously which is different from our approach where only one region can be focused at a time.
@@
Another family of models combines object-centric representations (Greff et al., 2019; Locatello et al., 2020) with relational inductive biases (Webb et al., 2021; Kerg et al., 2022; Webb et al., 2024b). The former decomposes an image into objects present in it while the latter constrains the processing to relations between objects instead of their specific properties. Webb et al. (2024a) used pre-trained slot attention (Locatello et al., 2020) to segment objects into a sequence of slots that contain disentangled representations of objects’ visual features and their positional information. Each pair of slots is then processed by a relational operator and the resulting sequence of relational representations is processed by the Transformer (Vaswani et al., 2017). The successor model, Slot-Abstractor (Mondal et al., 2024), processes the sequence of slots with a recently introduced Abstractor module (Altabaa et al., 2024) that proved to be especially effective in reasoning about relations between objects using their disentangled visual and positional features. Similar to our approach, there are two streams of information: one based on visual features and one based on positional information. However, the latter stream consists of positional embeddings that are learned during the pre-training process which may make the model more fragile to OOD inputs. In contrast, our approach uses the raw 2D locations of salient regions. Interestingly, in the domain of visually extended large language models (LLMs), Bhattacharyya et al. (2024) showed that forcing these models via surrogate tasks to collect relevant pieces of information before solving the actual task improves the final performance. While using a different approach compared to ours to achieve that, this work points to the potential of integrating our conceptual approach into LLMs.
@@ 4 Experiments
Our approach was evaluated on four visual reasoning datasets illustrated in Figure 3. We start with an evaluation on all 23 tasks of the SVRT dataset (Fleuret et al., 2011) to see how efficiently the model can learn spatial and similarity relations between simple shapes, see an example in Figure 3(a). Next, we test the model’s OOD capabilities to generalize learned relations to previously unseen shapes. Specifically, we use an extended version of the task #1 from the SVRT dataset (Puebla & Bowers, 2022), referred to as SVRT #1-OOD, see Figure 3(b). It contains test sets with OOD shapes coming from 13 different distributions. Additionally, we use the ART dataset (Webb et al., 2021), which contains more abstract relations, see Figure 3(c). Finally, we test the model’s potential to scale to more complex images by testing it on the CLEVR-ART dataset (Webb et al., 2024a), see Figure 3(d). It contains two tasks from ART but with more realistic images. Each dataset is described in detail in Appendix B. While all tasks described above consist of simple images, they are still challenging, as illustrated by the limited accuracy of the baselines described below.
@@ 5 Results
In all experiments, we report the performance of GAP combined with two downstream architectures, Transformer and Abstractor. The corresponding models are referred to as GAP-Transformer and GAP-Abstractor respectively. Details about hyper-parameters and training process are described in Appendix C. One important hyper-parameter is the type of glimpse sensor (multi-scale or log-polar) that we select for each of the four considered datasets. We report performance for different glimpse sensors in Appendix  D. Our code is publicly available at https://github.com/IBM/glimpse-based-active-perception.
Table 1: Test accuracy for SVRT evaluated for common dataset sizes of 1000 and 500 samples. The results obtained from the best trained models are averaged over SD, SR and over all (SD + SR) tasks. The standard deviation is calculated over accuracies for each individual task. Prior art results are listed in the top section whereas results obtained from our evaluations are listed in the two bottom sections. The highest and the second-highest mean accuracy in each column is marked in bold and underlined, respectively. The models appended by * contain a module pre-trained on the train sets from all SVRT tasks to segment the shapes.
@@
Figure 4: Sample efficiency evaluation for SVRT beyond common dataset sizes of 1000 and 500 samples. Average test accuracy on all 23 SVRT tasks depending on the size of the training dataset. Panel (a) compares our best model with prior art. Panel (b) compares our glimpse-based models with their ablated counterparts.
@@ 5.2 Out-of-distribution generalization of same-different relation
While the SVRT dataset allows us to see whether a model can learn visual relations from a limited amount of training data, it does not allow us to test the model’s ability to generalize those relations to OOD inputs. We evaluate the OOD generalization of a single same-different relation using the SVRT #1-OOD dataset (Appendix  B.2) where a model has to determine whether two shapes are the same or different up to translation. Importantly, the models are evaluated on test sets with OOD shapes coming from 13 different distributions (see Figure B.2 for examples). Table 5.2 shows the performance averaged over all OOD test sets and Figure 5 shows the performance on each of the test sets. Outperforming the prior art models, our model, GAP-Abstractor, performs quite consistently on all test sets being slightly less accurate on only two sets, ’straight_lines’ and ’rectangles’. Figure B.3 illustrates the reason why these two sets are more challenging.
@@ 5.4 Scaling potential for more complex visual scenes
We test our approach on two tasks from ART – relational-match-to-sample (RMTS) and identity rules (ID) – but with more realistic images using the CLEVR-ART dataset. In the RMTS task the model has to determine whether the objects in the top row are in the same “same/different” relation as the objects in the bottom row. In the ID task the model has to determine whether the bottom row contains objects that follow the same abstract pattern (ABA, ABB or AAA) as the objects in the top row (see Figure B.5). While the images from other datasets considered in this work are binary and contain only 2D objects, the CLEVR-ART images are colored and contain 3D objects with such real-world features as shades and reflections. Hence, the glimpsing behavior can be distracted by spurious salient locations such as edges of shades or regions of high reflection. Table 4 highlights that the GAP-Abstractor performs on par with the state-of-the-art model, Slot-Abstractor, being marginally better in one task and marginally worse in the other. However, Slot-Abstractor was pre-trained on all shapes from CLEVR-ART being thereby potentially exposed to objects from both train and test sets. This is in contrast to our models, which were not pre-trained and were exposed only to the objects from the train set.
@@
To compare our models with Slot-Abstractor on a more equal footing, we generated two additional OOD test sets shown in Figure 6. The new OOD-1 test set contains one novel type of shape – pyramid – on which the Slot-Abstractor was not pre-trained. Similarly, the new OOD-2 test set contains 2 such novel shapes: pyramids and toruses. The results are shown in Figure 7 demonstrating superior OOD generalization of our GAP-based model compared with Slot-Abstractor. In addition, we provide preliminary experiments with more realistic objects and more complex saliency map extraction in Appendix E.
@@ 6 Discussion
Our results suggest that factorizing an image into its complementary “what” and “where” contents plays an essential role. The low-dimensional “where” content (glimpse locations) is crucial since it does not contain any visual information and, in turn, allows to learn relations that are agnostic to specific visual details. To capitalize on that, it is important to process the “where” content explicitly, disentangling it from its “what” complement. We implemented this by using the recently introduced relational cross-attention mechanism (employed by the Abstractor downstream architecture) where the “what” content is used to compute the attention scores for the processing of the “where” content. In contrast, implicitly mixing the “what” and the “where” contents, as in the case of the Transformer downstream architecture, weakens the generalization capabilities. This can be seen by comparing the performance between GAP-Transformer and GAP-Abstractor. To further support this, we provide supplementary results in Appendix D.2 showing the importance of using both the “what” and the “where” contents instead of just one of them for task solving.
Another important aspect of our model is that it processes only the most salient image regions while ignoring the unimportant ones. The inferior performance of models where the downstream architectures receive full information (i.e. all patches that constitute the image) suggests that supplying a superfluous amount of information may distract the models hindering the learning process. Additional results provided in Appendix D.3 elucidate this further. Specifically, we show that it is insufficient to naively reduce the amount of information provided to the downstream architecture, by discarding uninformative image parts. Instead, it has to be ensured that the supplied information is well structured. In the context of GAP, it means that the glimpse contents have to be centered around salient image regions. For example, the image patches of the glimpse contents should contain objects’ edges in their central parts rather than somewhere in the periphery.
@@ 7 Conclusion
Acknowledgments
The authors would like to thank Wolfgang Maass for the helpful comments and fruitful discussions.
@@ B.2 SVRT #1-OOD
Figure B.3: Difficult cases of SVRT #1-OOD: The shown images are from the ’straight_lines’ and ’rectangles’ OOD test sets. The images contain different shapes that might be difficult to distinguish even for humans.
@@ Appendix D Ablations and additional results
GAP-Transformer multi-scale 76.4 ± 4.3 64.7 ± 5.6
@@
D.1 Performance differences between glimpse sensors
We evaluated two different glimpse sensors for each of the four datasets considered in this work, see Table D.1D.4 and Figure D.6. For the Abstractor-based downstream architecture, we observe comparable performance for both sensors in most cases. The most significant performance difference is observed for SVRT #1-OOD and two tasks from ART dataset (Table D-D.3) where the log-polar sensor achieves around 5% better accuracy. For the Transformer-based downstream architecture, the performance difference between two glimpse sensors averaged over all tasks is around 5%. However, there are tasks where the multi-scale sensor results in around 10% worse accuracy, see Table D.1, and 10% better accuracy, see Table D.3. Overall, based on the performance data, there is no clear quantitative trend favoring either of the two glimpse sensors. We ascribe this to the qualitative advantages and disadvantages of both sensors. In particular, the log-polar sensor benefits from the properties of the log-polar space where the rotation and scaling in the Cartesian space become translations, see e.g. (Javier Traver & Bernardino, 2010). However, the warping into log-polar space may make it difficult to capture information about image parts that are far from the glimpse locations. The multi-scale sensor, in contrast, captures the distant information more faithfully using downsampled image patches of different sizes. The disadvantage of this, however, is that the resulting glimpse content contains additional dimensions for the different scales.
@@ D.1 Performance differences between glimpse sensors
Dataset size SR tasks
@@
D.2 Evaluating importance of glimpse contents and glimpse locations
@@
D.3 Evaluating the role of processing only the relevant information
@@
Appendix E Experiment with more realistic objects and alternative saliency maps
We evaluated our model on more realistic objects. Additionally, we show how different saliency maps may improve the model’s performance in this case.
In particular, we used the objects from the Super-CLEVR dataset (Li et al., 2023), including cars, planes and motorbikes, to generate a new test set for the Identity Rules task from the CLEVR-ART dataset. The task is to determine whether three objects in the top row are arranged under the same abstract rule as the objects the bottom row. For example, objects “bus”, “car” and “bus” in one row are arranged under the abstract pattern ABA just as objects “car”, “bus” and “car. We considered four models trained on the original CLEVR-ART train set that consists only of simple cubes (Figure 6), and evaluated them on the new dataset, referred to as Super-CLEVR test set. The example images from the Super-CLEVR test set are shown in Figure E.8.
Simultaneously, we explored using different saliency maps. The first model is our GAP-Abstractor model with the original error neurons applied on the raw image to extract the saliency map, we refer to this model here as GAP-Abstractor (EN). The second model, referred to as GAP-Abstractor (CNN+EN), is a modification of GAP-Abstractor that extracts the saliency map by applying the error neurons on the feature maps computed by the first layer of ResNet-18 that was pre-trained on ImageNet. Computing the saliency map from the feature maps should make the glimpsing behavior less distracted by spurious salient locations, such as edges of shades or regions of high reflection. The third model, referred to as GAP-Abstractor (IK), employs an alternative way to extract the saliency maps using a more complex bio-plausible model proposed by Itti et al. (1998). The example saliency maps are shown in Figure E.8. Table E.8 shows the corresponding results, where Slot-Abstractor is the best performing prior-art model providing the baseline.
Figure E.8: Different saliency maps for images with more realistic objects. The images were generated using objects from the Super-CLEVR dataset for the ID task defined by the CLEVR-ART dataset.
Table E.8: Accuracy [%] on ID task from CLEVR-ART on the original test set and the OOD test with more realistic objects from Super-CLEVR. The accuracy results are averaged over 5 trained models. The highest and the second-highest mean accuracy in each column is marked in bold and underlined, respectively.
Model
original
test set
Super-CLEVR
test set
Slot-Abstractor (prior art) 91.61 ± 0.2 24.1 ± 0.8
GAP-Abstractor (EN) 93.4 ± 1.3 72.3 ± 2.3
GAP-Abstractor (CNN+EN) 92.7 ± 1.8 80.2 ± 1.9
GAP-Abstractor (IK) 90.2 ± 2.1 82.1 ± 2.2
As it can be seen, Slot-Abstractor fails completely on the unseen Super-CLEVR objects while the performance of our GAP-Abstractor does not degrade so severely. Moreover, we see that GAP-Abstractor (CNN+EN) performs better on Super-CLEVR test set while performing comparably on the original test set. Therefore, applying the error neurons on top of features maps rather than on top of the raw images results in a saliency map that can better steer the glimpsing behavior. Lastly, we see that GAP-Abstractor (IK) with an alternative bio-plausible way to extract the saliency map reaches the best performance on the Super-CLEVR test set while being slightly worse on the original test set. A possible reason for this difference is that this saliency map extractor was designed to handle real-world scenes that are of much higher visual complexity compared to simple shapes. Overall, these results indicate a possible direction of enhancing our approach for visual reasoning by integrating more powerful saliency maps.